# A Galactic Transformation—Understanding the SMC’s Structural and Kinematic Disequilibrium

### v1

@@ Abstract
software: This work made use of python, and its packages like numpy (vanderWalt2011; Harris2020), scipy (Virtanen2020) and matplotlib (Hunter2007).
@@
1 Introduction
Figure 1: Left panel: Placing the SMC on the Baryonic Tully-Fisher Relation (BTFR). The black dashed line denotes the BTFR fit taken from McGaugh2000, with the grey shaded region denoting the $1-\sigma$ uncertainty. The SMC’s total baryonic mass (stars ++ HI) is $(7.0\pm 0.6)\times 10^{8}$ M (Stanimirovic1999; Harris2004). Symbols denote the SMC’s inferred peak HI rotation velocity of $56\pm 5$ km s-1 (DiTeodoro2019, DT19) and the peak old star rotation velocity of <10<10 km s-1 (Zivick2021, Z21). The kinematics of the SMC’s HI and old stars are strongly discrepant. The orange circle denotes the SMC’s initial disk in the B12 simulation, which is consistent with BTFR by design. Right panel: the SMC’s photometric center $(12.80^{\circ},-72.83^{\circ})$ (Gonidakis2009) and the HI kinematic center $(16.26^{\circ},-72.42^{\circ})$ (Stanimirovic2004) are plotted over a background of SMC stars selected from the NN Optimal Gaia DR3 sample of Arranz2023. The two centers are separated by 121-2 kpc on-sky, which is a significant fraction of the SMC’s on-sky extent of $\approx 4-5$ kpc. Explaining the kinematic discrepancies between the old stars and HI (rotation peaks and centers) is the goal of this work.
@@ 1 Introduction
The SMC is the most massive satellite galaxy of the LMC, and is the second most massive satellite of the Milky Way (MW). Given the SMC’s proximity ($\sim 60$ kpc, Cioni2000), a wealth of astrometric, photometric, and spectroscopic data is available for this galaxy from both ground- and space-based observatories. As such, the SMC is a benchmark for studies of star formation, the interstellar medium (ISM), and dust physics in the low metallicity regime (e.g. RusselDopita1992; Gordon1998; Murray2024Scylla). Hence, the SMC is a standard calibrator used to study high-redshift galaxies (e.g. Kulkarni2021; Schouws2022; Joseph2023; Markov2023; Roman-Duval2025; Ocvirk2025). Further, with the advent of Gaia, the internal stellar kinematics of the SMC are now well constrained (Luri2021; Zivick2021; Dhanush2025), making the SMC a potential unique laboratory for dark matter (DM) physics.
@@
However, Gaia proper motions (PM) reveal that the SMC is in a high state of disequilibrium (Zivick2021; Luri2021). Hence, equilibrium assumptions cannot be used to describe the SMC’s ISM structure and kinematics, or the SMC’s DM distribution. Theoretical models of the SMC’s disequilibrium are urgently required—not only to interpret the SMC’s observational data, but also to assesss the reliability of the SMC as a calibrator for low-metallicity galaxies. In this work, we present a new theoretical framework to understand the origin of the SMC’s disequilibrium in structure and kinematics, in the context of its interaction history with the LMC.
The SMC’s disequilibrium is clearly illustrated by placing it on the Baryonic Tully-Fisher relation (BTFR, TF1977; McGaugh2000). The SMC has a gas-to-stellar mass ratio close to 1 (Stanimirovic1999; Harris2004; Bruns2005). In the local universe, dwarf galaxies with such high gas fractions generally possess rotation supported disks (e.g. The LITTLE THINGS survey – Hunter2012; Johnson2012; Johnson2015), see also Geha2006. Indeed, local gas-rich dwarfs satisfy the BTFR (e.g. Geha2006; Begum2008; McGaugh2021). This implies that the SMC should also satisfy the BTFR. In isolated disk galaxies where stellar kinematics are measurable, both stars and gas serve as tracers of the same underlying rotational field (Hinz2003; Williams2010). In Figure 1, the SMC is placed on the BTFR using the rotation amplitude derived from both stars and gas. The SMC’s old stars (age >> 1 Gyr) and gas (HI) are strongly discrepant in the BTFR (see also BekkiChiba2008; BekkiChiba2009)—illustrating the SMC’s gas and stars are not in equilibrium. The old stellar kinematics is highly dispersion dominated ($v_{\rm rot}/\sigma<0.6$, Harris2006), with a small rotation amplitude (<20<20 km s-1, Hardy1989; Harris2006; Dobbie2014; Zivick2018; Niederhofer2021; Zivick2021; Dhanush2025). However, HI surveys have revealed a velocity gradient of 60-100 km s-1 across the SMC’s $\sim 5^{\circ}$ extent. This has been interpreted as the signature of a rotating gas disk with a peak velocity of $\approx$ 50 km s-1 (e.g. Hindman1967; Stanimirovic2004; DiTeodoro2019; Pingel2022).
@@
The spatial distribution of the SMC’s old stellar populations and gas also points to the system being in a state of disequilibrium. The SMC has a large stellar line-of-sight (LoS) depth, which ranges from 5 kpc to 20 kpc depending on the measurement technique (Mathewson1986; Hatzidimitriou1993; Crowl2001; Subramanian2012; Nidever2013; Ripepi2017; Subramanian2017; Muraveva2018; Tatton2021; Dhanush2025). HI surveys show a bi-modal gas LoS velocity profile (Johnson1961; Hindman1967; Torres1987; Pingel2022), with two brightness temperature peaks separated by $\approx$50 km s-1 due to gas clumps 5–10 kpc apart along the LoS (Murray2024). Understanding the SMC’s complex gas distribution is crucial to properly interpret maps of dust extinction (Murray2024) and HI column densities (Pingel2022), which are ultimately used to derive star formation rates (Bolatto2011; Jameson2016). Further, understanding the large stellar LoS depth is important for converting PMs to 3D velocities. This work investigates the origin of the SMC’s large stellar LoS depth and the bi-modal LoS gas distribution.
@@
Unsurprisingly, the SMC’s center is ill-defined given the discrepant stellar and gas kinematics, and the complex 3-D structure. The photometric center derived from near Infra-Red (IR) star counts (Gonidakis2009) is separated from the HI kinematic center (Stanimirovic2004) by 121-2 kpc on-sky, as illustrated in right panel of Figure 1. For reference, the SMC is $\sim$4 kpc across on the sky, meaning a spatial difference of 1 - 2 kpc in stellar- and gas-derived centers must point to a high degree of disequilibrium. This study aims to provide a physical explanation for the separation between the SMC’s different centers, enabling us to identify the correct center for analyzing stellar and gas kinematics.
@@
The coincident star formation histories of the SMC and LMC over the last $\sim 4$ Gyr (Harris2009; Weisz2013; Massana2022) suggest that the Clouds have been bound for at least that amount of time. Further, the formation of the $150^{\circ}$ long gas stream (MS) (Mathewson1974; Braun2004; Nidever2010) and the SMC-LMC gas bridge (Kerr1957; Putman2003; Bruns2005) require the Clouds to complete multiple pericenters about each other (e.g. Murai1980; Besla2010; Pardy2018; Lucchini2021). There is increasing observational evidence that the most recent (100200100-200 Myr ago) SMC-LMC pericenter was a direct collision between the Clouds with an impact parameter of $\sim 2$ kpc (Olsen2011; Besla2016; Zivick2018; Zivick2019; Choi2022; Dhanush2024; Arranz2025; Rathore2025a; Rathore2025b).
@@
In this work, we use hydrodynamic simulations of the SMC-LMC-MW interaction history from Besla2012 to: (i) understand how a recent ($\sim 100$ Myr ago) SMC-LMC direct collision (impact parameter $\approx 2$ kpc) affects the SMC’s internal structure and kinematics; (ii) provide a means to interpret observations of the SMC’s stars and gas; and (iii) assess the validity of equilibrium assumptions for measuring the SMC’s DM content.
@@
This manuscript is organized as follows. Section 2 gives an overview of the B12 simulations and presents a coordinate system for analyzing the simulated SMC. Then, we study the effects of the SMC-LMC collision on the SMC’s stellar morphology (section 3.1), stellar kinematics (section 3.2), and gas kinematics (section 3.3). In section 4, we investigate the LoS distribution of the simulated SMC’s gas, understand the effect of the LMC ISM’s ram pressure on the SMC’s gas during the collision, discuss ideas to determine the SMC’s DM content without relying on equilibrium assumptions, compare our results to previous studies and assess the limitations of our analysis and prospects for the future. We conclude in section 5.
2 Methods
2.1 Hydrodynamic Simulations of the LMC-SMC-MW Interactions
@@
The initial LMC and SMC are modeled with live Hernquist DM halos (Hernquist1990), live exponential stellar disks and Smooth Particle Hydrodynamic (SPH) gas disks. The DM, stellar and gas masses of the LMC (SMC) were chosen to be $1.76\times 10^{11}\>(2\times 10^{10})$ M, $2.5\times 10^{9}\>(2.6\times 10^{8})$ M and $1.1\times 10^{9}\>(7.9\times 10^{8})$ M respectively. The scale radii of the LMC (SMC) stellar and gas disks were chosen to be 1.7 (1.1) kpc and 1.7 (3.3) kpc. The DM, stellar and gas mass resolution for the LMC (SMC) was $1.76\times 10^{6}\>(1.4\times 10^{6})$ M, 2500(2600)2500\>(2600) M and 3667(2633)3667\>(2633) M respectively. The MW was modeled as a static NFW (NFW1997) potential, with a total mass of $1.5\times 10^{12}$ M, Virial concentration of 12 and Viral radius of 300 kpc.
@@
The B12 simulation was performed using the Gadget-3 N-body SPH code (Springel2005). A sub-resolution multiphase model with an effective equation of state was used for the Interstellar medium (ISM) (Springel2003). Radiative cooling and star formation were implemented according to the Springel2003 prescription.
@@
Both Model 1 and Model 2 are in reasonable agreement with several external features associated with the Clouds. These features include: the observed SMC’s trailing gas stream (MS) (Mathewson1974; Braun2004; Nidever2010) and its stellar counterpart (Chandra2023; Zaritsky2025); the SMC-LMC gas bridge (Kerr1957; Putman2003; Bruns2005); and the LMC’s galactocentric position and velocity at present day as determined by Kallivayalil2006. We refer the reader to the B12 paper and Besla2013 for a detailed comparison between the aforementioned observed features and the simulation.
The SMC-LMC collision scenario in the Model 2 simulation is particularly successful at explaining the LMC’s observed internal features. These features include - the LMC’s outer disk morphology (Besla2016), disturbed kinematics of the LMC’s old (age >> 1 Gyr) stellar populations (Choi2022), the disturbed morphology and kinematics of the LMC’s bar (Rathore2025b; Rathore2025a), and the LMC’s gas distribution (Rathore2025b, hereafter R25). Further, the observed stellar PM vectors in the LMC-SMC bridge are also in reasonable agreement with Model 2 (Zivick2019). None of the above observations are explainable in Model 1, which does not invoke a collision.
@@
As with any simulation, B12 Model 2 is just one possibility of the SMC - LMC - MW interaction history, and the model has several limitations. We will discuss these limitations and how they affect our main conclusions in detail in section 4.5. We emphasize that our aim is not to exactly represent the observed SMC’s structure and kinematics with the B12 simulation. Rather, we want to use the B12 simulation to understand the physical mechanisms that operate when the SMC collides with the LMC, and provide a framework to interpret the SMC’s observations.
One limitation of Model 2, which is clear from Figure 2, is that the simulated SMC-LMC separation at present day ($\approx 10$ kpc) is inconsistent with the observed separation between the Clouds (20 kpc, Kallivayalil2006). Given this limitation, we will refrain from labeling the epochs in the simulation with respect to the present day. Rather, we will analyze the simulated SMC’s structure and kinematics as a function of time elapsed since the SMC-LMC collision (key times are marked in Figure 2). This approach will help us understand the rate at which the SMC’s disequilibrium grows post-collision.
@@
2.2 Centering and Alignment of the Simulated SMC
@@
We compute the SMC’s stellar density center with an iterative shrinking sphere algorithm. We start with a guess for the density center, which is the mass weighted average of the positions of all SMC stars. Then, we consider a spherical volume of radius 5 kpc centered at the guess density center. The mass weighted average position of stars that reside inside that spherical volume is evaluated. This gives a new estimate for the density center. We re-center the sphere at the new density center, shrink its radius by 30%, and re-compute the mass weighted average position of stars that reside inside the shrunken sphere. Convergence is established when the density center estimates between two successive iterations differ by less than 0.1 kpc, which corresponds to the softening length of the simulation. We experiment with different values of the hyper-parameters (like the starting radius of the sphere, shrinking fraction, convergence threshold), finding that the above adopted values give the best estimate of the density center (see Figure 3). Such an iterative approach is commonly used to calculate density centers for disturbed stellar and dark matter distributions in simulations and observations (Power2003; GC2019; Rathore2025a; Rathore2025b).
@@
The identification of the SMC’s stellar kinematic center post-collision is especially tricky. Computing the stellar kinematic center requires the azimuthal velocities, consequently requiring a plane of rotation. A standard practice in simulations is to use the total angular momentum vector of the stellar velocity field to identify the plane of rotation (e.g. Gomez2016; Gomez2017; Rathore2025b). But, computing the total angular momentum vector requires us to already know the kinematic center. For isolated disks, the stellar density center can be used as a good measure of the kinematic center. However, for highly disturbed kinematics, i.e. a post-collision system, the kinematic center cannot be apriori assumed to coincide with the density center. Hence, the kinematic center and the total angular momentum vector (or equivalently the plane of rotation) of the post-collision SMC cannot be determined independently of each other. Needed is an iterative approach that can jointly constrain both the kinematic center and the total angular momentum vector.
@@
3 Results
3.1 The Simulated SMC’s Stellar Structure

### v2

@@ Abstract
software: This work made use of python, and its packages like numpy (van der Walt et al., 2011; Harris et al., 2020), scipy (Virtanen et al., 2020) and matplotlib (Hunter, 2007).
@@
I Introduction
[…]
@@ I Introduction
The SMC is the most massive satellite galaxy of the LMC, and is the second most massive satellite of the Milky Way (MW). Given the SMC’s proximity ($\sim 60$ kpc, Cioni et al., 2000), a wealth of astrometric, photometric, and spectroscopic data is available for this galaxy from both ground- and space-based observatories. As such, the SMC is a benchmark for studies of star formation, the interstellar medium (ISM), and dust physics in the low metallicity regime (e.g. Russell and Dopita, 1992; Gordon and Clayton, 1998; Murray et al., 2024b). Hence, the SMC is a standard calibrator used to study high-redshift galaxies (e.g. Kulkarni et al., 2021; Schouws et al., 2022; Lewis et al., 2023; Markov et al., 2023; Roman-Duval et al., 2025; Ocvirk et al., 2025). Further, with the advent of Gaia, the internal stellar kinematics of the SMC are now well constrained (Gaia Collaboration et al., 2021; Zivick et al., 2021; Dhanush et al., 2025), making the SMC a potential unique laboratory for dark matter (DM) physics.
@@
However, Gaia proper motions (PM) reveal that the SMC is in a high state of disequilibrium (De Leo et al., 2020; Zivick et al., 2021; Gaia Collaboration et al., 2021). Hence, equilibrium assumptions (e.g. Ostriker et al., 2010) cannot be used to describe the SMC’s ISM structure and kinematics, or the SMC’s DM distribution. Theoretical models of the SMC’s disequilibrium are urgently required—not only to interpret the SMC’s observational data, but also to assesss the reliability of the SMC as a calibrator for low-metallicity galaxies. In this work, we present a new theoretical framework to understand the origin of the SMC’s disequilibrium in structure and kinematics, in the context of its interaction history with the LMC.
[…]
@@
The spatial distribution of the SMC’s old stellar populations and gas also points to the system being in a state of disequilibrium. The SMC has a large stellar line-of-sight (LoS) depth, which ranges from 5 kpc to 20 kpc depending on the measurement technique (Mathewson et al., 1986; Hatzidimitriou et al., 1993; Crowl et al., 2001; Subramanian and Subramaniam, 2012; Nidever et al., 2013; Ripepi et al., 2017; Subramanian et al., 2017; Muraveva et al., 2018; Tatton et al., 2021; Yanchulova Merica-Jones et al., 2021; Almeida et al., 2024; Dhanush et al., 2025; Oden et al., 2025). HI surveys show a bi-modal gas LoS velocity profile (Johnson, 1961; Hindman, 1967; Torres and Carranza, 1987; Pingel et al., 2022), with two brightness temperature peaks separated by $\approx$50 km s-1 due to gas clumps 5–10 kpc apart along the LoS (Murray et al., 2024a). Understanding the SMC’s complex gas distribution is crucial to properly interpret maps of dust extinction (Murray et al., 2024a) and HI column densities (Pingel et al., 2022), which are ultimately used to derive star formation rates (Bolatto et al., 2011; Jameson et al., 2016). Further, understanding the large stellar LoS depth is important for converting PMs to 3D velocities. This work investigates the origin of the SMC’s large stellar LoS depth and the bi-modal LoS gas distribution.
@@
Unsurprisingly, the SMC’s center is ill-defined given the discrepant stellar and gas kinematics, and the complex 3-D structure. The photometric center derived from near Infra-Red (IR) star counts (Gonidakis et al., 2009) is separated from the HI kinematic center (Stanimirović et al., 2004) by 121-2 kpc on-sky, as illustrated in right panel of Figure 1 (see also De Leo et al. 2020). For reference, the SMC is $\sim$4 kpc across on the sky, meaning a spatial difference of 1 - 2 kpc in stellar- and gas-derived centers must point to a high degree of disequilibrium. This study aims to provide a physical explanation for the separation between the SMC’s different centers, enabling us to identify the correct center for analyzing stellar and gas kinematics.
@@
The coincident star formation histories of the SMC and LMC over the last $\sim 4$ Gyr (Noël et al., 2007, 2009; Harris and Zaritsky, 2009; Weisz et al., 2013; Massana et al., 2022; Burhenne et al., 2026) suggest that the Clouds have been bound for at least that amount of time. Further, the formation of the $150^{\circ}$ long gas stream (MS) (Mathewson et al., 1974; Braun and Thilker, 2004; Nidever et al., 2010) and the SMC-LMC gas bridge (Kerr and Hindman, 1957; Putman et al., 2003; Brüns et al., 2005) require the Clouds to complete multiple pericenters about each other (e.g. Murai and Fujimoto, 1980; Besla et al., 2010; Pardy et al., 2018; Lucchini et al., 2021). There is increasing observational evidence that the most recent (100200100-200 Myr ago) SMC-LMC pericenter was a direct collision between the Clouds with an impact parameter of $\sim 2$ kpc (Olsen et al., 2011; Besla et al., 2016; Zivick et al., 2018, 2019; Choi et al., 2022; Dhanush et al., 2024; Jiménez-Arranz and Roca-Fàbrega, 2025; Rathore et al., 2025b, a).
@@
In this work, we use hydrodynamic simulations of the SMC-LMC-MW interaction history from Besla et al. (hereafter B12, 2012) to: (i) understand how a recent ($\sim 100$ Myr ago) SMC-LMC direct collision (impact parameter $\approx 2$ kpc) affects the SMC’s internal structure and kinematics; (ii) provide a means to interpret observations of the SMC’s stars and gas; and (iii) assess the validity of equilibrium assumptions for measuring the SMC’s DM content.
@@
This manuscript is organized as follows. Section II gives an overview of the B12 simulations and presents a coordinate system for analyzing the simulated SMC. Then, we study the effects of the SMC-LMC collision on the SMC’s stellar morphology (section III.1), stellar kinematics (section III.2), and gas kinematics (section III.3). In section IV, we investigate the LoS distribution of the simulated SMC’s gas, understand the effect of the LMC ISM’s ram pressure on the SMC’s gas during the collision, discuss ideas to determine the SMC’s DM content without relying on equilibrium assumptions, compare our results to previous studies and assess the limitations of our analysis and prospects for the future. We conclude in section V.
II Methods
II.1 Hydrodynamic Simulations of the LMC-SMC-MW Interactions
@@
The initial LMC and SMC are modeled with live Hernquist DM halos (Hernquist, 1990), live exponential stellar disks and Smooth Particle Hydrodynamic (SPH) gas disks. The DM, stellar and gas masses of the LMC (SMC) were chosen to be $1.76\times 10^{11}\>(2\times 10^{10})$ M, $2.5\times 10^{9}\>(2.6\times 10^{8})$ M and $1.1\times 10^{9}\>(7.9\times 10^{8})$ M respectively. The scale radii of the LMC (SMC) stellar and gas disks were chosen to be 1.7 (1.1) kpc and 1.7 (3.3) kpc. The scale heights of the LMC (SMC) stellar and gas disks were chosen to be 0.34 (0.11) kpc and 0.34 (0.33) kpc respectively. These chosen values of the scale radii and scale height are consistent with isolated dIrr galaxies in the local universe (Swaters et al., 2002; Connors et al., 2004; Kreckel et al., 2011). The gravitational softening length for the gas and stars was 0.1 kpc.
The DM, stellar and gas mass resolution for the LMC (SMC) was $1.76\times 10^{6}\>(1.4\times 10^{6})$ M, 2500(2600)2500\>(2600) M and 3667(2633)3667\>(2633) M respectively. The MW was modeled as a static NFW (Navarro et al., 1997) potential, with a total mass of $1.5\times 10^{12}$ M, Virial concentration of 12 and Viral radius of 300 kpc.
For the full table of simulation parameter choices and their justification, we refer the reader to Table 1 of B12.
@@
The B12 simulation was performed using the Gadget-3 N-body SPH code (Springel, 2005). A sub-resolution multiphase model with an effective equation of state was used for the Interstellar medium (ISM) (Springel and Hernquist, 2003). Radiative cooling and star formation were implemented according to the Springel and Hernquist (2003) prescription.
Note that ram pressure from the Clouds’ ISM might play an important role in shaping the gas morphology and kinematics of the Clouds. However, ram pressure is known to be underestimated in standard SPH (e.g. Agertz et al., 2007). We will discuss the role of ram pressure in some detail in section IV.1, and provide an analytical means to understand the effect of ram pressure on the Clouds.
@@
Both Model 1 and Model 2 are in reasonable agreement with several external features associated with the Clouds. These features include: the observed SMC’s trailing gas stream (MS) (Mathewson et al., 1974; Braun and Thilker, 2004; Nidever et al., 2010) and its stellar counterpart (Chandra et al., 2023; Zaritsky et al., 2025); the SMC-LMC gas bridge (Kerr and Hindman, 1957; Putman et al., 2003; Brüns et al., 2005); and the LMC’s galactocentric position and velocity at present day as determined by Kallivayalil et al. (2006b). We refer the reader to the B12 paper and Besla et al. (2013) for a detailed comparison between the aforementioned observed features and the simulation.
The SMC-LMC collision scenario in the Model 2 simulation is particularly successful at explaining the LMC’s observed internal features. These features include - the LMC’s outer disk morphology (Besla et al., 2016), disturbed kinematics of the LMC’s old (age >> 1 Gyr) stellar populations (Choi et al., 2022), the disturbed morphology and kinematics of the LMC’s bar (Rathore et al., 2025a, b), and the LMC’s gas distribution (Rathore et al., 2025a, hereafter R25). Further, the observed stellar PM vectors in the LMC-SMC bridge are also in reasonable agreement with Model 2 (Zivick et al., 2019). None of the above observations are explainable in Model 1, which does not invoke a collision.
@@
As with any simulation, B12 Model 2 is just one possibility of the SMC - LMC - MW interaction history, and the model has several limitations. We will discuss these limitations and how they affect our main conclusions in detail in section IV.5. We emphasize that our aim is not to exactly represent the observed SMC’s structure and kinematics with the B12 simulation. Rather, we want to use the B12 simulation to understand the physical mechanisms that operate when the SMC collides with the LMC, and provide a framework to interpret the SMC’s observations.
One limitation of Model 2, which is clear from Figure 2, is that the simulated SMC-LMC separation at present day ($\approx 10$ kpc) is inconsistent with the observed separation between the Clouds (20 kpc, Kallivayalil et al., 2006b). Given this limitation, we will refrain from labeling the epochs in the simulation with respect to the present day. Rather, we will analyze the simulated SMC’s structure and kinematics as a function of time elapsed since the SMC-LMC collision (key times are marked in Figure 2). This approach will help us understand the rate at which the SMC’s disequilibrium grows post-collision.
The B12 simulation does qualitatively match the relative locations of the LMC, SMC and MW with respect to each other. More specifically, the simulated SMC is located south of the LMC at present day, and is at a larger Galactocentric distance as compared to the Galactocentric distance of the LMC. Quantitatively, the SMC’s Galactocentric $x$ and $z$ positions are off by $\approx 10$ and $\approx 5$ kpc respectively when compared to observations. The simulated SMC’s Galactocentric velocities at present day are in reasonable agreement with observations from Kallivayalil et al. (2006b, a). As B12 mention, a more accurate orbital solution for the SMC can be achieved by fine tuning the simulation initial conditions. However, such fine tuning is unlikely to change the physical mechanisms that govern the LMC-SMC collision. For a more detailed comparison of the simulated phase space of the Clouds to the observations, we refer the reader to Table 2 of B12.
@@
II.2 Centering and Alignment of the Simulated SMC
@@
We compute the SMC’s stellar density center with an iterative shrinking sphere algorithm. We start with a guess for the density center, which is the mass weighted average of the positions of all SMC stars. Then, we consider a spherical volume of radius 5 kpc centered at the guess density center. The mass weighted average position of stars that reside inside that spherical volume is evaluated. This gives a new estimate for the density center. We re-center the sphere at the new density center, shrink its radius by 30%, and re-compute the mass weighted average position of stars that reside inside the shrunken sphere. Convergence is established when the density center estimates between two successive iterations differ by less than 0.1 kpc, which corresponds to the softening length of the simulation. We experiment with different values of the hyper-parameters (like the starting radius of the sphere, shrinking fraction, convergence threshold), finding that the above adopted values give the best estimate of the density center (see Figure 3). Such an iterative approach is commonly used to calculate density centers for disturbed stellar and dark matter distributions in simulations and observations (Power et al., 2003; Garavito-Camargo et al., 2019; Rathore et al., 2025b, a).
@@
The identification of the SMC’s stellar kinematic center post-collision is especially tricky. Computing the stellar kinematic center requires the azimuthal velocities, consequently requiring a plane of rotation. A standard practice in simulations is to use the total angular momentum vector of the stellar velocity field to identify the plane of rotation (e.g. Gómez et al., 2016, 2017; Rathore et al., 2025a). But, computing the total angular momentum vector requires us to already know the kinematic center. For isolated disks, the stellar density center can be used as a good measure of the kinematic center. However, for highly disturbed kinematics, i.e. a post-collision system, the kinematic center cannot be apriori assumed to coincide with the density center. Hence, the kinematic center and the total angular momentum vector (or equivalently the plane of rotation) of the post-collision SMC cannot be determined independently of each other. Needed is an iterative approach that can jointly constrain both the kinematic center and the total angular momentum vector.
@@
III Results
III.1 The Simulated SMC’s Stellar Structure